OpenStudy (anonymous):

the sum of two consecutive integers is 284. write an equation and solve to find the two integers

OpenStudy (anonymous):

Did you write the problem correctly

OpenStudy (anonymous):

idk how to write the problem

OpenStudy (anonymous):

if they are consecutive, then they have a +1 relationship: (x) + (x+1) = 284

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

However, you will not get integers because integers are any positive or negative whole number. With the information given, I believe, you will get rational answers.

OpenStudy (anonymous):

2x+1=284 2x= 283 X= 141.5. Not an integer X+1=142.5. Not an integer